Given the coordinate transformation
u[sub]1[/sub] = xy
2u[sub]2[/sub] = x[sup]2[/sup] + y[sup]2[/sup]
u[sub]3[/sub] = z,
Determine if the coordinate system is orthogonal.
1
Expert's answer
2011-06-22T08:00:14-0400
The coordinate system is ortogonal if the metric tensor is diagonal.
Orthogonal coordinates never have off-diagonal terms in their metric tensor.
metric tensor Gij=Summ(L=1 to N) (duL/dxi+duL/dxj)
(x y z)=(x1 x2 x3)
So we must check if the tensor is diagonal
G11 = y2+1+0 = 1+y2
G12 = xy+xy+0 = 2xy = G21
G13 = 0+0+0=0 = G31
G22 = 0+y2+0 = y2
G33 = 0+0+1 = 1
So we can say that given coordinate system isn't ortogonal